OpenStudy (anonymous):

In 2006 the median price of an existing home in some country was approximately $ 225,400. In 2001 the median price was $ 153,000. Let Y be the median price of the existing home in the year x, where x=0 represents 2001. 1. write a linear equation that models the median existing home price in terms of the year x/ y= ( write it in slope intercept form) 2.Use the equationto predict the median existing home price for the year 2010. Y=$ 3. Interpret the slope of the equation found in part 1.

OpenStudy (agent47):

Ok so suppose you start your x axis from the year 2000, or x=0, corresponds to the year 2000. That would mean in 2001: x=1 and y=153,000 In 2006: x=6 and y=225,400

OpenStudy (agent47):

So you have two points: (1;153,000) and (6;225,400) M=(y2-y1)/(x2-x1)=(225,400-153,000)/(6-1)=72400/5=14480

OpenStudy (agent47):

Equation of the line is: y-153,000=14,480(x-1); y=14,480x-14,480+153,000=14,480x+167,480

OpenStudy (agent47):

So part 1: y=14,480x+167,480 Part 2: Plug in x=10 for 2010, and see what y is: y=14,480*10+167,480=144,800+167,480=$312,280

OpenStudy (agent47):

OMG MISTAKES!

OpenStudy (agent47):

lol I realized I made a dumb mistake when I was explaining part 3... pheeww

OpenStudy (anonymous):

its ok u make this so simple to learn

OpenStudy (agent47):

y=14,480x-14,480+153,000 y=14,480x+138520

OpenStudy (agent47):

In that step, instead of subtracting 14,480, I added.

OpenStudy (agent47):

So part 1: y=14,480x+138520 part 2: y=14,480*10+138,520 = 283320

OpenStudy (agent47):

part 3: Slope is rise over run, which, in our case is the amount by which the price increases by every year. If it were 2, our price would double every year, if it were 4, it would quadruple every year, etc.. And, they're not asking you for it, but +138,520 in that equation shows that the initial price was $138,520.

OpenStudy (anonymous):

so what would part threes answer be?